OpenStudy (anonymous):

The ratio of two numbers is 7 to 5. The sum of the numbers is 24. What are the numbers? (Points : 4) 12 and 12 14 and 10 21 and 15 18 and 6

OpenStudy (freckles):

say the numbers are x and y ratio of x to y = ratio of 7 to 5 means x/y=7/5 and the sum of x and y is 24 means x+y=24 so you have x/y=7/5 and x+y=24

OpenStudy (freckles):

solve the first equation for x by multiplying both sides by y

OpenStudy (freckles):

then sub that into the other equation